OpenStudy (anonymous):

DEFG is a rectangle. DF = 5x-5 and EG=x+11. Find the value of x and the length of each diagonal.

OpenStudy (anonymous):

4x = 16 x= 4 l = 5*4 -5 =15 l2 = 4+11 =15